A dozen eggs cost $0.88 in January 1980 and $2.11

in January 2015. The average wage for production

workers was $7.58 per hour in January 1980 and

$19.64 in January 2015.

a. By what percentage did the price of eggs rise?

b. By what percentage did the wage rise?

c. In each year, how many minutes did a worker

have to work to earn enough to buy a

dozen eggs?

d. Did workers’ purchasing power in terms of eggs

rise or fall?


1
Expert's answer
2022-01-23T15:39:52-0500

a)"\frac{2.11}{0.88}\times100=239.77"

price increase by 139.77 %


b)"\frac{19.67}{7.58}\times100=259.10"

increase by 159.10 %


c) 1980:

"\frac{0.88\times60}{7.58}=6.97"


2015:

"\frac{2.11\times60}{19.64}=6.45"


d)Workers’ purchasing power in terms of eggs is rise.